# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v2

@@ 1 Introduction
The explanations for this tension fall into four categories. First, star formation might be intrinsically more efficient at Cosmic Dawn (Dekel et al., 2023). Second, a low mass-to-light ratio at Cosmic Dawn might cause models calibrated to the lower-redshift Universe to overestimate stellar masses. These explanations include the presence of active galactic nuclei (AGN) (Wang et al., 2024), a top-heavy stellar initial mass function (IMF) (Sharda & Krumholz, 2022; Cameron et al., 2023), a lack of dust along the line of sight (Mason et al., 2023; Ferrara et al., 2023), and differences in dust properties. Third, observations could be biased by bursty star formation at Cosmic Dawn (Sun et al., 2023; Shen et al., 2023). In this scenario, the bright end of the UVLF is dominated by less intrinsically bright galaxies which have undergone a recent burst. Finally, a modified $\Lambda$-CDM cosmology can produce a greater abundance of massive haloes at Cosmic Dawn. These explanations include “Early Dark Energy” (Klypin et al., 2021; Shen et al., 2024) or massive primordial black hole seeds (Liu & Bromm, 2022).
@@
At high redshift, massive haloes become increasingly rare. Typical comoving volumes of JWST galaxy surveys are $\sim 10^{5}$ to $10^{6}~{}{\rm cMpc^{3}}$. For example, in the Cosmic Evolution Early Release Science (CEERS Finkelstein et al., 2024, ERS-1345), the effective survey volume at UV magnitude $\leq-20$ is $\simeq 160000~{}{\rm cMpc^{3}}$. Thus, the haloes of observed galaxies should have comoving number densities $\gtrsim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc})^{-3}$. In Figure 1, we plot the comoving number density of dark matter haloes as a function of halo mass and redshift, assuming cosmological parameters from the Plank 2018 results (Planck Collaboration et al., 2020). We estimate the number density by integrating the Press-Schechter halo mass function (Press & Schechter, 1974) using the Python package COLOSSUS (Diemer, 2018). We also plot an estimate for the typical baryon surface density of the galaxy hosted by the halo. The calculation is described in Appendix A.
Haloes at redshifts $z\sim 9$ and masses $\sim 10^{11}~{}M_{\odot}$ (marked by the red star) are interesting objects because they have number densities $\gtrsim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc})^{-3}$ and they are expected to host galaxies with high baryon surface densities $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$. For comparison, a Milky-Way-like galaxy at $z=0$ with halo mass $10^{12}~{}M_{\odot}$ has the baryon surface density $\sim 3\times 10^{2}~{}M_{\odot}/{\rm pc^{2}}$. We expect the difference in surface densities to be even greater for the gas component, because at low redshift, massive galaxies tend to be quenched with most of their baryons in stars. We can also estimate the typical Hydrogen number density by assuming a disc scale height (App. A). For a scale height $(H/R)_{\rm gal}=0.1$ and 50% of the baryon mass in gas, the number density is $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$.
Figure 1: The comoving number density of dark matter haloes as a function of halo mass and redshift. A simple estimate for baryon surface density (App. A) is used to plot lines of constant baryon surface density in red. The most massive galaxies at Cosmic Dawn seen by JWST likely have comoving number densities $\sim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc^{3}})^{-3}$. Our simulation (red star) probes galaxies with similar number densities. The high baryon surface densities $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ and number densities $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ create a unique environment for star formation which is not realized in the lower redshift Universe.
@@
The physical conditions in MDGs differ from present day galaxies in other ways too. The rapid accretion of cold streams in massive galaxies at high redshift is expected to seed strong turbulence (Dekel et al., 2009; Ginzburg et al., 2022). This is empirically supported by the high velocity dispersions observed in galaxies at $z\gtrsim 6$ (de Graaff et al., 2024b). Turbulent pressure supports molecular clouds against gravitational collapse, suppressing local star formation. Turbulence also reduces the clustering of supernovae (SNe), making SN feedback less effective (Gentry et al., 2017). In addition to turbulence, the low metal content of the ISM and the higher CMB temperature at Cosmic Dawn may suppress star formation by making cooling less efficient.
Galaxy formation during the epoch of reionization is often studied with big box cosmological simulations such as MassiveBlack (Di Matteo et al., 2012), CROC (Gnedin, 2014), Eagle (Schaye et al., 2015; Crain et al., 2015), BlueTides (Feng et al., 2016), Aurora (Pawlik et al., 2017), Illustris-Tng (Naiman et al., 2018; Nelson et al., 2018; Marinacci et al., 2018; Springel et al., 2018; Pillepich et al., 2018), Sphinx (Rosdahl et al., 2018), SIMBA (Davé et al., 2019), Cosmic Dawn II (Ocvirk et al., 2020), and Thesan (Kannan et al., 2022). However, it is challenging to simultaneously resolve the structure of the ISM at scales $\sim 10~{}{\rm pc}$ and use a large enough volume $\simeq(100h^{-1}{\rm cMpc})^{3}$ to sample rare massive haloes.
@@ 2.1 RAMSES
We assume an ideal gas equation of state (EOS) appropriate for a rarefied plasma with $\gamma=5/3$. The turbulent kinetic energy (TKE) $\varepsilon_{\rm turb}$ and a refinement mask $M$ are advected passively with the flow. The trajectories of dark matter and star cluster particles are computed using a particle-mesh solver. The Poisson equation is solved using a multi-grid method (Guillet & Teyssier, 2011) with Dirichlet boundary conditions at refinement level boundaries.
@@
Criterion (i) ensures that only cells which exchange gas with the initial zoom-in region are refined. Criterion (iv) ensures that we satisfy the Truelove et al. (1997) criterion in cells that are not maximally refined, preventing artificial fragmentation. Cells that are maximally refined and Jeans unstable will rapidly form stars before they have time to fragment. Therefore, we do not require an artificial pressure floor as used in previous works (Robertson & Kravtsov, 2008; Agertz et al., 2009).
@@ 2.2 Subgrid turbulence model
The bulk components are defined as the mass-weighted volume-average fields, also known as Favre-averaged fields. The specific TKE is defined as
$\varepsilon_{\rm turb}=\frac{1}{2}\overline{\rho v^{\prime\prime 2}}=\frac{1}{2}\overline{\rho}\sigma_{\rm 3D}^{2}=\frac{3}{2}\overline{\rho}\sigma_{\rm 1D}^{2}$ (4)
@@
$\partialderivative{t}\varepsilon_{\rm turb}+\partialderivative{x_{j}}(\varepsilon_{\rm turb}\widetilde{v}_{j})+P_{\rm turb}\partialderivative{\widetilde{v}_{j}}{x_{j}}=\mathcal{C}_{\rm turb}-\mathcal{D}_{\rm turb}$ (5)
where $P_{\rm turb}=(2/3)\varepsilon_{\rm turb}=\overline{\rho}\sigma_{\rm 1D}^{2}$ is the turbulent pressure and $\mathcal{C}_{\rm turb}$ and $\mathcal{D}_{\rm turb}$ are creation and destruction source terms respectively.
@@ 3.2 Galaxy structure
The gas and star mass is organized into dense clumps of size $\sim 100~{}{\rm pc}$, although the clump size may be limited by our effective resolution $\simeq 10~{}{\rm pc}$. Nonetheless, high-redshift disc galaxies show similar structures in observation (Elmegreen et al., 2007; Guo et al., 2012, 2018; Soto et al., 2017) and simulation (Noguchi, 1998; Inoue et al., 2016; Mandelker et al., 2017; Mandelker et al., 2024), strengthening the analogy. The recent simulation work of (Nakazato et al., 2024) find that similar clump structures at high redshifts $z\gtrsim 5$ might be detectable by JWST due to their brightness in rest-frame optical emission lines.
@@ 3.7 Variable turbulence forcing parameter
We leave a more rigorous analysis of self-consistent turbulence forcing to future work which simulations the forcing parameter more carefully. Mandelker et al. (2024) describe a more rigorous approach to implement a self-consistent forcing parameter which involves decomposing the shear tensor into three components representing compression, solid body rotation and shear flow. Their method will be described in an upcoming paper (Ginzburg et. al., in prep.).
@@ 4.4 Turbulence
In real galaxies, the relationship between the forcing of small-scale turbulence and large-scale galactic dynamics is complex. varTurb provides a first insight into how this could play out. Solenoidal forcing modes in the turbulence may be driven by the differential rotation within a disc, suppressing star formation in discs. Meanwhile, compressive forcing modes may be driven by SN explosions, galaxy-galaxy mergers (Renaud et al., 2014), or accretion (Mandelker et al., 2024), enhancing star formation in those environments.
@@ 4.9.2 Stellar masses and SFRs
The crude comparison in Figure 16 is reasonable because the stellar masses and SFRs inferred from observations are already highly model-dependent. We are only suggesting that our simulated galaxies are in the ballpark of stellar masses and SFRs of the MDGs seen by JWST. Future work could make this comparison more quantitative using abundance matching and accounting for the effects of cosmic variance (Kragh Jespersen et al., 2024), but a more detailed approach may not be justified until the observational data are better understood.
More observational evidence for MDGs are found at Cosmic Noon ($z\sim 3$), near the peak of the cosmic SFR density, where a population of massive quiescent galaxies (MQGs) have emerged (Glazebrook et al., 2017; Valentino et al., 2020; Antwi-Danso et al., 2023; Carnall et al., 2023, 2024; de Graaff et al., 2024a; Glazebrook et al., 2024). These galaxies typically have large stellar masses $M_{*}\gtrsim 10^{10}~{}M_{\odot}$ and low specific SFRs $\lesssim 0.2$. Recently, Weibel et al. (2024) reported the discovery of a MQG at $z=7.3$ confirmed spectroscopically with JWST NIRSpec as part of the Cycle 2 program of RUBIES (GO#4233) with a stellar mass $M_{*}\sim 10^{10}~{}M_{\odot}$.
Detailed modeling of MQG spectra suggests that MQGs must have formed the bulk of their stellar populations within the first billion years of cosmic history in short, extreme starbursts lasting hundreds of ${\rm Myr}$ at most Weibel et al. (e.g. 2024). This is qualitatively consistent with our simulated MDGs, hinting that MDGs could be the progenitors of MQGs. Future work should extend run simulations out to lower redshift to investigate this possibility. Cosmological simulations tend to underproduce MQGs at high redshift (Alcalde Pampliega et al., 2019; Merlin et al., 2019), so they have been understudied in the theoretical galaxy formation community.
@@ 5 Conclusions and Future Work
The interplay between photoionization and SN feedback in our simulations suggests that early feedback is important in MDGs, even when it cannot suppress star formation directly. In future simulations, we will model early feedback in more detail using RAMSES-RT, including explicit models for photoionization, radiation pressure, and stellar winds. These models should account for the effect of strong turbulence, which can decrease the effective opacity (Sec. LABEL:sec:turb_deatt). Using RAMSES-RT will also enable a more self-consistent treatment of the extragalactic radiation field.

### v3

@@ 1 Introduction
The explanations for this tension fall into four categories. First, star formation might be intrinsically more efficient at Cosmic Dawn (Dekel et al., 2023). Second, a low mass-to-light ratio at Cosmic Dawn might cause models calibrated to the lower-redshift universe to overestimate stellar masses. These explanations include the presence of active galactic nuclei (AGN) (Wang et al., 2024), a top-heavy stellar initial mass function (IMF) (Sharda & Krumholz, 2022; Cameron et al., 2024), a lack of dust along the line of sight (Mason et al., 2023; Ferrara et al., 2023), and differences in dust properties. Third, observations could be biased by bursty star formation at Cosmic Dawn (Sun et al., 2023; Shen et al., 2023). In this scenario, the bright end of the UVLF is dominated by less intrinsically bright galaxies which have undergone a recent burst. Finally, a modified $\Lambda$-CDM cosmology can produce a greater abundance of massive haloes at Cosmic Dawn. These explanations include “Early Dark Energy” (Klypin et al., 2021; Shen et al., 2024) or massive primordial black hole seeds (Liu & Bromm, 2022).
@@
At high redshift, massive haloes become increasingly rare. Typical comoving volumes of JWST galaxy surveys are $\sim 10^{5}$ to $10^{6}~{}{\rm cMpc^{3}}$. For example, in the Cosmic Evolution Early Release Science (CEERS Finkelstein et al., 2024, ERS-1345), the effective survey volume at UV magnitude $\leq-20$ is $\simeq 160000~{}{\rm cMpc^{3}}$. Thus, the haloes of observed galaxies should have comoving number densities $\gtrsim 10^{-6}~{}{\rm cMpc}^{-3}$. In Figure 1, we plot the comoving number density of dark matter haloes as a function of halo mass and redshift, assuming cosmological parameters from the Plank 2018 results (Planck Collaboration et al., 2020). We estimate the number density by integrating the Press-Schechter halo mass function (Press & Schechter, 1974) using the Python package COLOSSUS (Diemer, 2018). We also plot an estimate for the typical baryon surface density of the galaxy hosted by the halo. The calculation is described in Appendix A.
Haloes at redshifts $z\sim 9$ and masses $\sim 10^{11}~{}M_{\odot}$ (marked by the red star) are interesting objects because they have number densities $\sim 2\times 10^{-5}~{}{\rm cMpc^{-3}}$, so a small sample of such objects should exist within current surveys. Simultaneously, they are expected to host galaxies with high baryon surface densities $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$. For comparison, a Milky-Way-like galaxy at $z=0$ with halo mass $10^{12}~{}M_{\odot}$ has number density $\sim 2\times 10^{-3}~{}{\rm cMpc^{-3}}$ and baryon surface density $\sim 3\times 10^{2}~{}M_{\odot}/{\rm pc^{2}}$. We expect the difference in surface densities to be even greater for the gas component, because at low redshift, massive galaxies tend to be quenched with most of their baryons in stars. We can also estimate the typical Hydrogen number density by assuming a disc scale height (App. A). For a scale height $(H/R)_{\rm gal}=0.1$ and 50% of the baryon mass in gas, the number density is $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$.
Figure 1: The comoving number density of dark matter haloes as a function of halo mass and redshift. A simple estimate for baryon surface density (App. A) is used to plot lines of constant baryon surface density in red. The most massive galaxies at Cosmic Dawn seen by JWST likely have comoving number densities $\sim 10^{-6}~{}{\rm cMpc}^{-3}$. Our simulation (red star) probes galaxies with similar number densities. The high baryon surface densities $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ and number densities $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ create a unique environment for star formation which is not realized in the lower redshift Universe.
@@
The physical conditions in MDGs differ from present day galaxies in other ways too. The rapid accretion of cold streams in massive galaxies at high redshift is expected to seed strong turbulence (Dekel et al., 2009; Ginzburg et al., 2022). This is empirically supported by the high velocity dispersions observed in galaxies at $z\gtrsim 6$ (de Graaff et al., 2024). Turbulent pressure supports molecular clouds against gravitational collapse, suppressing local star formation. Turbulence also reduces the clustering of supernovae (SNe), making SN feedback less effective (Gentry et al., 2017). In addition to turbulence, the low metal content of the ISM and the higher CMB temperature at Cosmic Dawn may suppress star formation by making cooling less efficient.
Galaxy formation during the epoch of reionization is often studied with big box cosmological simulations such as MassiveBlack (Di Matteo et al., 2012), CROC (Gnedin, 2014), Eagle (Schaye et al., 2015; Crain et al., 2015), BlueTides (Feng et al., 2016), Aurora (Pawlik et al., 2017), Illustris-Tng (Naiman et al., 2018; Nelson et al., 2018; Marinacci et al., 2018; Springel et al., 2018; Pillepich et al., 2018), Sphinx (Rosdahl et al., 2018), SIMBA (Davé et al., 2019), Cosmic Dawn II (Ocvirk et al., 2020), and Thesan (Kannan et al., 2022). However, it is challenging to simultaneously resolve the structure of the ISM at scales $\sim 10~{}{\rm pc}$ and use a large enough volume $\simeq(100h_{0}^{-1}{\rm cMpc})^{3}$ to sample rare massive haloes.
@@ 2.1 RAMSES
We assume an ideal gas equation of state (EOS) appropriate for a rarefied plasma with $\gamma=5/3$. The specific turbulent kinetic energy (TKE) $\varepsilon_{\rm turb}$ and a refinement mask $M$ are advected passively with the flow. The trajectories of dark matter and star cluster particles are computed using a particle-mesh solver. The Poisson equation is solved using a multi-grid method (Guillet & Teyssier, 2011) with Dirichlet boundary conditions at refinement level boundaries.
@@
where the Jeans length is $\lambda_{\rm J}=(\pi c_{\rm s}^{2}/G\rho)^{1/2}$ and $c_{\rm s}=\sqrt{\gamma P/\rho}$ is the sound speed.
Criterion (i) ensures that only cells which exchange gas with the initial zoom-in region are refined. Criterion (iv) ensures that we satisfy the Truelove et al. (1997) criterion $\Delta x/\lambda_{\rm J}<0.25$ in cells that are not maximally refined, preventing artificial fragmentation.
Artificial fragmentation can still occur when maximally refined cells violate the Truelove criterion. However, these cells only persist for one freefall time due to rapid star formation, limiting the development of artificial fragmentation. In our fiducial simulation at $z=9$, only 6% of maximally refined cells violate the Truelove criterion. In future simulations, we may use a pressure floor to eliminate artificial fragmentation altogether, following previous works (Robertson & Kravtsov, 2008; Agertz et al., 2009).
@@ 2.2 Subgrid turbulence model
The bulk components are defined as the mass-weighted volume-average fields, also known as Favre-averaged fields. The TKE is defined as
$e_{\rm turb}=\frac{1}{2}\overline{\rho v^{\prime\prime 2}}=\frac{1}{2}\overline{\rho}\sigma_{\rm 3D}^{2}=\frac{3}{2}\overline{\rho}\sigma_{\rm 1D}^{2}$ (4)
@@
$\partialderivative{t}e_{\rm turb}+\partialderivative{x_{j}}(e_{\rm turb}\widetilde{v}_{j})+P_{\rm turb}\partialderivative{\widetilde{v}_{j}}{x_{j}}=\mathcal{C}_{\rm turb}-\mathcal{D}_{\rm turb}$ (5)
where $P_{\rm turb}=(2/3)e_{\rm turb}=\overline{\rho}\sigma_{\rm 1D}^{2}$ is the turbulent pressure and $\mathcal{C}_{\rm turb}$ and $\mathcal{D}_{\rm turb}$ are creation and destruction source terms respectively.
@@ 3.2 Galaxy structure
The gas and star mass is organized into dense clumps of size $\sim 100~{}{\rm pc}$, although the clump size may be limited by our effective resolution $\simeq 10~{}{\rm pc}$. Nonetheless, high-redshift disc galaxies show similar structures in observation (Elmegreen et al., 2007; Guo et al., 2012, 2018; Soto et al., 2017) and simulation (Noguchi, 1998; Inoue et al., 2016; Mandelker et al., 2017; Mandelker et al., 2025), strengthening the analogy. The recent simulation work of (Nakazato et al., 2024) find that similar clump structures at high redshifts $z\gtrsim 5$ might be detectable by JWST due to their brightness in rest-frame optical emission lines.
@@ 3.7 Variable turbulence forcing parameter
We leave a more rigorous analysis of self-consistent turbulence forcing to future work which simulates the forcing parameter more carefully. Mandelker et al. (2025) describe a more rigorous approach to implement a self-consistent forcing parameter which involves decomposing the shear tensor into three components representing compression, solid body rotation and shear flow. Their method will be described in an upcoming paper (Ginzburg et. al., in prep.).
@@ 4.4 Turbulence
In real galaxies, the relationship between the forcing of small-scale turbulence and large-scale galactic dynamics is complex. varTurb provides a first insight into how this could play out. Solenoidal forcing modes in the turbulence may be driven by the differential rotation within a disc, suppressing star formation in discs. Meanwhile, compressive forcing modes may be driven by SN explosions, galaxy-galaxy mergers (Renaud et al., 2014), or accretion (Mandelker et al., 2025), enhancing star formation in those environments.
@@ 4.9.2 Stellar masses and SFRs
The crude comparison in Figure 16 is reasonable because the stellar masses and SFRs inferred from observations are already highly model-dependent. We are only suggesting that our simulated galaxies are in the ballpark of stellar masses and SFRs of the MDGs seen by JWST. Future work could make this comparison more quantitative using abundance matching and accounting for the effects of cosmic variance (Jespersen et al., 2025), but a more detailed approach may not be justified until the observational data are better understood.
More observational evidence for MDGs are found at Cosmic Noon ($z\sim 3$), near the peak of the cosmic SFR density, where a population of massive quiescent galaxies (MQGs) have emerged (Glazebrook et al., 2017; Valentino et al., 2020; Antwi-Danso et al., 2025; Carnall et al., 2023, 2024; de Graaff et al., 2025; Glazebrook et al., 2024). These galaxies typically have large stellar masses $M_{*}\gtrsim 10^{10}~{}M_{\odot}$ and low specific SFRs $\lesssim 0.2$. Recently, Weibel et al. (2025) reported the discovery of a MQG at $z=7.3$ confirmed spectroscopically with JWST NIRSpec as part of the Cycle 2 program of RUBIES (GO#4233) with a stellar mass $M_{*}\sim 10^{10}~{}M_{\odot}$.
Detailed modeling of MQG spectra suggests that MQGs must have formed the bulk of their stellar populations within the first billion years of cosmic history in short, extreme starbursts lasting hundreds of ${\rm Myr}$ at most Weibel et al. (e.g. 2025). This is qualitatively consistent with our simulated MDGs, hinting that MDGs could be the progenitors of MQGs. Future work should extend run simulations out to lower redshift to investigate this possibility. Cosmological simulations tend to underproduce MQGs at high redshift (Alcalde Pampliega et al., 2019; Merlin et al., 2019), so they have been understudied in the theoretical galaxy formation community.
@@ 5 Conclusions and Future Work
The interplay between photoionization and SN feedback in our simulations suggests that early feedback is important in MDGs, even when it cannot suppress star formation directly. In future simulations, we will model early feedback in more detail using RAMSES-RT, including explicit models for photoionization, radiation pressure, and stellar winds. These models should account for the effect of strong turbulence, which can decrease the effective opacity (App. G). Using RAMSES-RT will also enable a more self-consistent treatment of the extragalactic radiation field.
@@ Acknowledgements
This research made use of SciPy (Virtanen et al., 2020), NumPy (Harris et al., 2020), and Matplotlib (Hunter, 2007).